# Global small data weak solutions of 2-D semilinear wave equations with scale-invariant damping, III

### v1

@@ Abstract
For the $2$-D semilinear wave equation with scale-invariant damping $\square u+\frac{\mu}{t}\partial_{t}u=|u|^{p}$, where $t\geq 1$, $\mu>0$ and $p>1$, it is conjectured that the global small data weak solution $u$ exists when $p>p_{s}(2+\mu)=\frac{\mu+3+\sqrt{\mu^{2}+14\mu+17}}{2(\mu+1)}$ for $0<\mu\leq 2$ and $p>p_{f}(2)=2$ for $\mu\geq 2$. In our previous papers, the global small solution $u$ has been obtained for $p>p_{s}(2+\mu)$ and $20<\mu<$ 2 but $\mu\not=1$. In the present paper, by the vector field method together with the delicate analysis on the Bessel functions, we will show the global existence of small solution $u$ for $p>2$ and $\mu\geq 2$. In forthcoming paper, for $\mu=1$ and $p>p_{s}(2+\mu)=p_{s}(3)=1+\sqrt{2}$, the global solution $u$ is also obtained. Therefore, this open question has been solved completely.
@@ 1 Introduction
In our forthcoming paper [12], (A1) with $\mu=1$ and $p>p_{s}(2+\mu)=1+\sqrt{2}$ will be shown by applying the Bessel function tools and deriving some suitable time-decay or spacetime-decay estimates of weak solution $u$ as in [1], [11] and [19]. Collecting the results in [19] and [11] (for $20<\mu<$ 2 but $\mu\not=1$, $p>p_{s}(2+\mu)$), [12] (for $\mu=1$, $p>p_{s}(2+\mu)=1+\sqrt{2}$) together with Theorem 1.1 in the paper (for $32<\mu<$ 3 and $p\geq 2$) , [3, Theorem 2] (for $\mu=2$ and $p\geq 2$) and [2, Theorem 2] (for $\mu\geq 3$ and $p\geq 2$), then the Open question (A) has been solved completely.
@@
Therefore, it follows from (1.3) that $\{Z^{I}\phi\}_{|I|\leq N_{0}}$ consist of a closed nonlinear wave system and then $\|Z^{I}\partial\phi\|_{L_{x}^{2}(\mathbb{R}^{n})}$ can be obtained by the standard energy estimate. By the crucial Klainerman-Sobolev inequality
@@
Note that for 2-D equation $\square u+\frac{2}{t}\,\partial_{t}u=|u|^{p}$, by the transformation $u=tv$, then the resulting semilinear wave equation $\square v=t^{-(p-1)}|v|^{p}$ can be obtained. By the standard vector field method and the energy estimate as in [21], the authors in Theorem 2 of [3] show that the small data solution $u\in C([1,\infty),H^{2}(\mathbb{R}^{2}))\cap C^{1}([1,\infty),H^{1}(\mathbb{R}^{2}))\cap C^{2}([1,\infty),L^{2}(\mathbb{R}^{2}))$ exists globally. However, for the general 2-D equation $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ with $32<\mu<$ 3, it is difficult to find a suitable transformation of the unknown function $u$ such that a related semilinear wave equation is derived. Moreover, the commutator $[\square+\frac{\mu}{t}\,\partial_{t},Z]$ can not be expressed an efficient linear combination of $\square+\frac{\mu}{t}\,\partial_{t}$ and $Z$. Indeed, letting $w_{1}=\partial_{t}u$ and $w_{2}=L_{1}u$, then one has from $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ that
@@
$\square w_{2}+\frac{\mu}{t}\,\partial_{t}w_{2}-\frac{\mu x_{1}}{t^{2}}w_{1}=L_{1}(|u|^{p}).$ (1.6)
This means that the linear parts in the nonlinear equations (1.5) and (1.6) become more complicated Euler-Poisson-Darboux-Klein-Gordon’s type models. To overcome this essential difficulty, we will take the following measures.
@@ 3 Estimates of solution to 2-D homogeneous equation $\square v+\frac{\mu}{t}\,\partial_{t}v=0$
$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{3}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho-1}=(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|,$ (3.59)
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{2}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho}=(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|.$
@@
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim t^{-1}|\xi|.$
@@ 4 Estimates of solution to 2-D inhomogeneous equation $\square w+\frac{\mu}{t}\,\partial_{t}w=F$
On the other hand, it follows from the Klainerman-Sobolev inequality in Chapter 6 of [13] that

### v2

@@ Abstract
For the $2$-D semilinear wave equation with scale-invariant damping $\square u+\frac{\mu}{t}\partial_{t}u=|u|^{p}$, where $t\geq 1$, $\mu>0$ and $p>1$, it is conjectured that the global small data weak solution $u$ exists when $p>p_{s}(2+\mu)=\frac{\mu+3+\sqrt{\mu^{2}+14\mu+17}}{2(\mu+1)}$ for $0<\mu\leq 2$ and $p>p_{f}(2)=2$ for $\mu\geq 2$. In our previous papers, the global small solution $u$ has been obtained for $p>p_{s}(2+\mu)$ and $20<\mu<$ 2 but $\mu\not=1$. In the present paper, by the vector field method together with the delicate analysis on the Bessel functions, we will show the global existence of small solution $u$ for $p>2$ and $\mu>2$. In forthcoming paper, for $\mu=1$ and $p>p_{s}(2+\mu)=p_{s}(3)=1+\sqrt{2}$, the global solution $u$ is also obtained. Therefore, collecting our series of conclusions together with partial results from others, this open question has been solved completely.
@@ 1 Introduction
In our forthcoming paper [12], (A1) with $\mu=1$ and $p>p_{s}(2+\mu)=1+\sqrt{2}$ will be shown by applying the Bessel function tools and deriving some suitable time-decay or spacetime-decay estimates of weak solution $u$ as in [1], [11] and [19]. Collecting the results in [19] and [11] (for $20<\mu<$ 2 but $\mu\not=1$, $p>p_{s}(2+\mu)$), [12] (for $\mu=1$, $p>p_{s}(2+\mu)=1+\sqrt{2}$) together with Theorem 1.1 in the paper (for $32<\mu<$ 3 and $p>2$) , [3, Theorem 2] (for $\mu=2$ and $p>2$) and [2, Theorem 2] (for $\mu\geq 3$ and $p>2$), then the Open question (A) has been solved completely.
@@
Therefore, it follows from (1.3) that $\{Z^{I}\phi\}_{|I|\leq N_{0}}$ consists of a closed nonlinear wave system and then $\|Z^{I}\partial\phi\|_{L_{x}^{2}(\mathbb{R}^{n})}$ can be obtained by the standard energy estimate. By the crucial Klainerman-Sobolev inequality
@@
Note that for 2-D equation $\square u+\frac{2}{t}\,\partial_{t}u=|u|^{p}$, by the transformation $u=tv$, then the resulting semilinear wave equation $\square v=t^{-(p-1)}|v|^{p}$ can be obtained. By the standard vector field method and the energy estimate as in [21], the authors in Theorem 2 of [3] show that the small data solution $u\in C([1,\infty),H^{2}(\mathbb{R}^{2}))\cap C^{1}([1,\infty),H^{1}(\mathbb{R}^{2}))\cap C^{2}([1,\infty),L^{2}(\mathbb{R}^{2}))$ exists globally. However, for the general 2-D equation $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ with $32<\mu<$ 3, it is difficult to find a suitable transformation of the unknown function $u$ such that a related semilinear wave equation is derived. Moreover, the commutator $[\square+\frac{\mu}{t}\,\partial_{t},Z]$ can not be expressed as an efficient linear combination of $\square+\frac{\mu}{t}\,\partial_{t}$ and $Z$. Indeed, letting $w_{1}=\partial_{t}u$ and $w_{2}=L_{1}u$, then one has from $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ that
@@
$\square w_{2}+\frac{\mu}{t}\,\partial_{t}w_{2}-\frac{\mu}{t}\partial_{1}u-\frac{\mu x_{1}}{t^{2}}w_{1}=L_{1}(|u|^{p}).$ (1.6)
This means that the linear parts in the nonlinear equations (1.5) and (1.6) become more complicated than the Euler-Poisson-Darboux operator together with the Klein-Gordon type’s operator. To overcome this essential difficulty, we will take the following measures.
@@ 3 Estimates of solution to 2-D homogeneous equation $\square v+\frac{\mu}{t}\,\partial_{t}v=0$
$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{3}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho-1}\leq(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|,$ (3.59)
$\left|\partial_{t}^{2}\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{2}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho}\leq(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|.$
@@
$\left|\partial_{t}^{2}\Psi_{1}(t,1,\xi)\right|\lesssim t^{-1}|\xi|.$
@@ 4 Estimates of solution to 2-D inhomogeneous equation $\square w+\frac{\mu}{t}\,\partial_{t}w=F$
On the other hand, it follows from the Klainerman-Sobolev inequality in [16] that